Trail Snax Corp. mixes raisins that cost $5.00 per kg with peanuts that cost $3.80 a kg. How many kilograms of raisins should be mixed with 10 kg of peanuts to obtain a mixture worth $4.00 per kilogram?
Answer by mananth(16946)

--------price ---------------- quantity
peanuts 3.8 ---------------- 10 kg
Raisins 5 ---------------- x kg
Mixture 4 ---------------- 10 + x kg

3.8*10+5x=4(10+ x)
38+5x=40+4x
5x+-4x=40-38
x=2
2 kg of Raisins
Peanuts = 8 kg